OpenStudy (anonymous):

A missile is launched from the ground. After 1 second, the missile is 103 feet in the air. After 2 seconds, the missile is 192 feet in the air. Using quadratic functions, find the height, in feet, of the missile after 5 seconds in the air.

OpenStudy (anonymous):

9,521 in total, decreasing in number as sites were overrun until October 1944.

OpenStudy (anonymous):

that didn't answer my question -_-

OpenStudy (anonymous):

Oh sorry -_-

OpenStudy (anonymous):

The problem can be solved because we're given 3 points which defines a parabola (similar to the other problem you posted racergirl) The 3 points are (0,0), (1,103), and (2,193) You set y=ax^2+bx+c and then you plug in for x for each of these points You know that c=0 right off the bat because when x=0, y=0 as well (a and b are both multiplied by zero when x=0) Then you have these two equations: a (1^2) + b(1)= 103 2^2 (a) + 2b=193 You solve the systems of equations (2 linear equations allows you to remove one of the variables) You get a=-7 and b=110 Then you plug in for x=5 to get a height of 375 ft

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem! did that make sense?

OpenStudy (anonymous):

kinda yeah.